Prove (cosAś1+sinA)+(1+sinAścosA)=2ścosA
Answer by Alan3354(69443)

Prove (cosAś1+sinA)+(1+sinAścosA)=2ścosA
-----
Prove cosA/(1+sinA)+(1+sinA)/cosA = 2/cosA
cos^2/(1+sin) + (1+sin) = 2
cos^2 + (1+sin)^2 = 2 + 2sin
1-sin^2 + 1 + 2sin + sin^2 = 2 + 2sin
2 + 2sin = 2 + 2sin
===========================
If anyone can post an example where working on one side only makes a difference, I would like to see it.